Use Gaussian Elimination only to solve

3x - 2y + z = 13

-2x + y + 4z = 11

x + 4y - 5z = -31
Answer #1

- 2 1 3x - 2y + z = 13 -2x + y + 42 = 1 7+49-52 = -31 - GA9- 3 2 4 Augmented matrix = (^12] I 3 - 2 -2 ( 4 11 14 -51 Swap Row